system of equations/elimination method
0.3x - 0.2y = 4
0.2x + 0.5y = 1

Answer by stanbon(75887)

system of equations/elimination method
0.3x - 0.2y = 4
0.2x + 0.5y = 1
--------------------
Get rid of the decimals:
3x - 2y = 40
2x + 5y = 10
--------------------
Multiply 1st Eq. by 2 and 2nd equation by 3 to get:
6x - 4y = 80
6x + 15y = 30
-----------------
Subtract 1st from 2nd to get:
19y = -50
y = -50/19
--------------------
Substitute into 2x+5y=10 to solve for "x":
2x + 5(-50/19) = 10
2x - 250/19 = 190/19
2x = 440/19
x = -220/19
------------------
Cheers,
Stan H.